# Deflection angle in the strong deflection limit: a perspective from local geometrical invariants and matter distributions

### v1

@@ VI Summary and discussion
In this paper, we have proposed a novel formulation of the deflection angle in the strong deflection limit for static, spherically symmetric spacetimes. By introducing an expansion parameter based on the areal radius—which has clear geometric significance—and isolating the logarithmically divergent part of the integral following the approach of Ref. [5], we derived the strong field limit coefficients that represent the divergence rate and the constant shift correction. Although these coefficients were previously expressed in a metric-dependent and coordinate-dependent manner, we have now related them to the second derivative of the effective potential—which governs photon dynamics near the photon sphere—and to the quasi-local mass. Furthermore, by relating these quantities to the tetrad components of the Einstein tensor, we have demonstrated that the strong field limit coefficients depend on local, coordinate-invariant curvatures. Since these curvatures are connected to various matter distribution quantities via the Einstein equations, we have ultimately succeeded in clarifying how the logarithmic divergence of the deflection angle in the strong deflection limit depends on matter field quantities. Note that since our results are expressed in terms of the Einstein tensor, they are, in essence, independent of the underlying gravitational theory.
@@
Our new formula clearly demonstrates that the logarithmic divergence rate, $\bar{a}$, is determined by the product of the square of the areal radius, $R_{\mathrm{m}}$, of the photon sphere and the sum of the local values of specific tetrad components of the Einstein tensor (i.e., $G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}})$. This result indicates that this quantity is the fundamental factor characterizing the strong deflection phenomenon occurring at the photon sphere. Within the framework of general relativity, the Einstein equations ensure that these local curvature quantities correspond to the local values of the energy density and the tangential pressure (i.e., $\rho_{\mathrm{m}}$ and $\Pi_{\mathrm{m}}$), implying that $\bar{a}$ depends solely on $R_{\mathrm{m}}^{2}\left(\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}\right)$. Therefore, by precisely measuring the logarithmic divergence rate of the deflection angle, we can directly determine $G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}}$, or, assuming general relativity, $\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}$. Moreover, the constant shift correction coefficient, $\bar{b}$, also reflects local curvature and quasi-local mass information; thus, a comprehensive evaluation of both coefficients is expected to yield deep insights into the internal structure and local physical conditions of compact objects.
@@
Our formulation clearly demonstrates why the special value $\bar{a}=1$ is ubiquitously observed across diverse spacetime models. In particular, at the photon sphere, when the matter fields satisfy $\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}=0$, the expression for $\bar{a}$ depends solely on $R_{\mathrm{m}}^{2}\left(\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}\right)$, inevitably resulting in $\bar{a}=1$. In every instance where $\bar{a}=1$ is observed, this condition is invariably satisfied. This insight crucially resolves the long-standing puzzle regarding the behavior of the logarithmic divergence rate.
@@
Furthermore, using our results, we show that within the eikonal approximation, the QNM frequencies can be expressed in terms of the local geometric and material properties near the photon sphere. The real part of the QNM frequencies is related to the orbital frequency at the photon sphere, while the imaginary part is determined by the Lyapunov exponent of the unstable photon circular orbit. In particular, we obtain
@@
where $n$ is the overtone number, $l$ is the angular momentum of the perturbation, and $\Omega_{\mathrm{c}}=1/b_{\mathrm{c}}$ denotes the angular coordinate velocity of the unstable photon circular orbit. The Lyapunov exponent is given by $\lambda_{\mathrm{L}}=b_{\mathrm{c}}^{-1}\sqrt{-V^{\prime\prime}_{\mathrm{m}}/2}$, which is related to $\bar{a}$ as [26]
@@
Thus, the QNM frequencies are ultimately determined by the same local matter field (or local curvature) quantities that govern the deflection angle. Therefore, precise measurements of QNM frequencies—such as those from gravitational wave observations—provide an independent means to probe the local curvature and matter distribution near the photon sphere, thereby reinforcing the connection between strong gravitational lensing and the dynamical properties of compact objects.
Our formulation is a universal, coordinate-independent framework that applies to a broader range of spacetime models than conventional methods, promising numerous future applications. Moreover, its independence from any specific gravitational theory enables its application to solutions beyond general relativity. Consequently, further investigation is needed to elucidate how the strong field limit coefficients depend on matter fields within modified gravitational theories, and our formulation is expected to serve as a foundation for clarifying their physical significance. While our current study is based on static, spherically symmetric spacetimes, extending the formulation to less symmetric spacetimes represents an important direction for future research. In addition, applying our approach to the strong deflection of massive particles offers another promising avenue for further investigation (for perturbations of the stable circular orbit of massive particles, see Ref. [38]). Furthermore, since a relationship between QNM frequencies and local curvature and matter fields is suggested, combining high-precision measurements of quasinormal modes from gravitational wave observations with analyses based on our formulation is expected to set the stage for a novel approach that yields detailed insights into the internal structure and local physical conditions of compact objects in extreme regimes.
@@ Acknowledgements.
This work was supported by JSPS KAKENHI Grant Nos. JP22K03611, JP23KK0048, JP24H00183.

### v2

@@ VI Summary and discussion
In this paper, we have proposed a novel formulation of the deflection angle in the strong deflection limit for static, spherically symmetric spacetimes. By introducing an expansion parameter based on the areal radius—which has clear geometric significance—and isolating the logarithmically divergent part of the integral, we have derived the strong field limit coefficients that characterize the divergence rate and the constant offset correction. While previous studies expressed these coefficients in a metric- and coordinate-dependent manner, our formulation relates them to the second derivative of the effective potential, which governs photon dynamics near the photon sphere, and to the photon sphere radius and the quasi-local mass. Furthermore, by relating these quantities to the tetrad components of the Einstein tensor, we have demonstrated that the strong field limit coefficients depend on local, coordinate-invariant curvatures. Since these curvatures are connected to matter distribution quantities via the Einstein equations, our results clarify how the logarithmic divergence of the deflection angle in the strong deflection limit is determined by matter field quantities. Note that since our results are expressed in terms of the Einstein tensor, they are, in essence, independent of the underlying gravitational theory.
@@
Our new formula demonstrates that the logarithmic divergence rate, $\bar{a}$, is determined by the product of the photon sphere’s areal radius squared, $R_{\mathrm{m}}^{2}$, and the sum of the specific tetrad components of the Einstein tensor, $G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}})$, evaluated at the photon sphere. This result indicates that $R_{\mathrm{m}}^{2}(G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}})$ is the fundamental factor characterizing the strong deflection phenomenon. Within the framework of general relativity, the Einstein equations ensure that these local curvature quantities correspond to the local energy density and tangential pressure (i.e., $\rho_{\mathrm{m}}$ and $\Pi_{\mathrm{m}}$), implying that $\bar{a}$ depends solely on $R_{\mathrm{m}}^{2}\left(\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}\right)$. Therefore, by precisely measuring the logarithmic divergence rate of the deflection angle, we can directly determine $G_{(0)(0)}^{\mathrm{m}}+G_{(2)(2)}^{\mathrm{m}}$, or, assuming general relativity, $\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}$. Moreover, the constant offset correction, $\bar{b}$, also encodes information about the local curvature and the quasi-local mass; thus, a comprehensive evaluation of both coefficients is expected to yield deep insights into the internal structure and local physical conditions of compact objects.
@@
Our formulation clearly explains why the special value $\bar{a}=1$ is ubiquitously observed across diverse spacetime models. In particular, at the photon sphere, if the matter fields satisfy $\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}=0$, then $R_{\mathrm{m}}^{2}\left(\rho_{\mathrm{m}}+\Pi_{\mathrm{m}}\right)$ vanishes, immediately implying $\bar{a}=1$. This insight crucially resolves the long-standing puzzle regarding the behavior of the logarithmic divergence rate.
@@
Furthermore, using our results, we show that within the eikonal approximation, the QNM frequencies can be expressed in terms of the local geometric and material quantities at the photon sphere. The real part of the QNM frequencies is related to the orbital frequency at the photon sphere, while the imaginary part is determined by the Lyapunov exponent of the unstable photon circular orbit. In particular, we obtain
@@
where $n$ is the overtone number, $l$ is the angular momentum of the perturbation, and $\Omega_{\mathrm{c}}=1/b_{\mathrm{c}}$ denotes the angular coordinate velocity of the unstable photon circular orbit. The Lyapunov exponent is given by $\lambda_{\mathrm{L}}=b_{\mathrm{c}}^{-1}\sqrt{-V^{\prime\prime}_{\mathrm{m}}/2}$, which is related to $\bar{a}$ as [28, 29]
@@
Thus, the QNM frequencies are ultimately determined by the same local matter field (or local curvature) quantities that govern the deflection angle. Therefore, precise measurements of QNM frequencies—such as those from gravitational wave observations—provide an independent means to probe the local curvature and matter distribution near the photon sphere, thereby reinforcing the connection between strong gravitational lensing and the dynamical properties of compact objects. Furthermore, our findings may also shed light on a universal upper bound on chaos in thermal quantum field theory [42, 43], which provides an inequality between the Lyapunov exponent and the surface gravity of the horizon, a relation whose generalization to the photon sphere has recently been discussed [44].
Our formulation is a universal, coordinate-independent framework that applies to a broader range of spacetime models than conventional approaches, promising numerous future applications. Moreover, its independence from any specific gravitational theory enables its application to solutions beyond general relativity. Consequently, further investigations are required to clarify how the strong field limit coefficients depend on matter fields in modified gravitational theories, and our formulation is expected to serve as a foundation for clarifying their physical significance. Although the present study focuses on static, spherically symmetric spacetimes, extending the formulation to less symmetric cases remains an important direction for future research. In addition, applying our formulation to the strong deflection of massive particles offers another promising avenue for further investigation (for perturbations of the stable circular orbit of massive particles, see Ref. [41]). Furthermore, since a relationship between QNM frequencies and local curvature and matter fields is suggested, combining high-precision measurements of quasinormal modes from gravitational wave observations with analyses based on our formulation is expected to set the stage for a novel approach that yields detailed insights into the internal structure and local physical conditions of compact objects in extreme regimes.
@@ Acknowledgements.
The author gratefully acknowledges the useful discussions with Ken-ich Nakao. This work was supported in part by JSPS KAKENHI Grant Nos. JP22K03611, JP23KK0048, and JP24H00183.